OpenStudy (anonymous):

the measure of angle J is half of its complement plus 12degrees

OpenStudy (anonymous):

complements add to 90 degrees. So if the measure of angle J is x, the measure of its complement is 90 - x. so: x = (90 - x)/2 + 12 x = 45 - x/2 + 12 3x/2 = 57 3x = 114 x = 38